OpenStudy (anonymous):

Find an equation for a line with a slope of 2 that has x-intercept of -3. Show your work then graph it. *someone help me please .

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

@YoungSWAHH are you sure that's "x-intercept" not "y-intercept"?

OpenStudy (anonymous):

Since it gives you the x intercept is -3 you know the coordinate for it is (-3,0) so you can use point slope form to figure out the equation. \[y-y_1=m(x-x_1)\]m=slope=2\[y_1=0\]\[x_1=-3\]